A population of kittens has black and white alleles that are codominant and the population consists of 14 black cats, 20 gray cats, and 30 white cats. calculate allele and genotype frequencies from this population and use to solve the main problem. Let B be the allele for black and b be the allele for white.
Main problem:An experiment is done with a population of 170 individuals from each of these genotypes. the survivor rates are: 150BB, 35Bb, 35bb
- relative fitness of the three genotypes
- avg fitness
- genotype frequencies in next generation based on fitness
Answer #1

Here BB = Black = 14

Bb = Grey = 20

bb = white = 30

Total cats = 64

1. For genotype frequency we use formula : Individual of specific genotype/ Total genotype:

for BB = 14/64 = 0.22

for Bb= 20/64 = 0.32

for bb = 30/64 = 0.46

1. For allele frequency we use formuala : total Individuals of specfic allele/ Total allele in population:

Here BB = 14 = 28 allele of B

Bb = 20 = 20 allele of B and 20 allele of b

bb = 30 = 60 allele of b

For B allel frequency = total Individuals of B allele/ Total allele in population = 48/128 = 0.37

For b allel frequency = total Individuals of b allele/ Total allele in population = 80/128 = 0.63

2. A. For relative fitness

Calculate the Relative Fitness (w) of each genotype by dividing each genotype'ssurvival and/or reproductive rate by the highest survival and/or reproductive rate among the 3 genotypes.

so for BB = 150/150 =1

for Bb 35/150 = 0.23

for bb 35/150 = 0.23

B. for average fitness

Take the Hardy-Weinberg equation (p2     2pq     q2) and multiply each term (the frequency of each genotype that is B and b) by the fitness of that genotype. Add those up and you get the mean fitness, w.

So for BB = p2 * 1 = 0.372 * 1 =0.13

for Bb = 2pq * 0.23 = 2 * 0.37 * 0.63 * 0.23 = 0.10

for bb = q2 * 0.23 = 0.632 * 0.23 = 0.091

Total average fitness = 0.13+0.10+0.091 = 0.32

C. genotype frequencies in next-generation based on fitness

The frequencies are going to be same since rand0m mating is going to happen adn the genotype frequency will be same in the next generation.